# Explain the Black Box for the Sake of Science: Revisiting the Scientific Method in the Era of Generative Artificial Intelligence

### v1

@@ Abstract
The framing of the problem is as follows. We want our AI approach to predict heatwaves in the Indochina peninsula 1 day ahead of time, using the prior 5 days. The associated learning task for our AI framework (based of a transformer neural network) is to correctly classify whether 1 day in the future corresponds or not to the onset day of a heatwave. We train our AI framework on time series data corresponding to 4 different variables that are relevant to heatwaves for the 5 days prior to the onset (or no onset) of a heatwave, and on a geographical domain larger than the Indochina peninsula. The training dataset is composed of 365 samples, the validation set of 77 samples, and the testing set also of 77 samples. The classification results show an overall accuracy on the test set of 88.3% (68/77), with a true positive (heatwave event) rate of 76.5% (13/18) and a true negative (no event) rate of 91.6% (55/59). From these results, we may argue that the machine learnt some ‘principles’ related to heatwaves in Indochina; especially if we focus on the 13 heatwave cases correctly classified in the test set.
@@
We would like now to understand what data the machine used to learn these ‘principles’. To this end, we adopt post-hoc interpretability, and in particular the methods and evaluation framework presented in turbe2023evaluation , to guarantee that the relevance maps (i.e., the machine view) are correctly identified – this satisfies the accuracy pillar set out in the discussion above. We then compute the average machine view across all samples that are correctly classified, for the 4 variables considered, and for the 5 days prior to a heatwave striking that were used as inputs. This point guarantees that the second and third pillar are also satisfied, that is: the relevance maps show reproducibility across correctly classified samples, and the variables considered are understandable to domain experts.
@@
What can we learn from the ‘machine view’ just retrieved? We can look at what domain experts understand of heatwaves in Indochina. To this end, we found only one relevant reference luo2018synoptic , that shows how temperature anomalies in South China seem to be systematic precursors of Indochina’s heatwaves, and constitutes the human view on this problem (as reported in Fig. 3). Remarkably, this human view and the machine view are convergent, at least partially – they deem the same temperature data as important for heatwaves in Indochina in the days preceding these events, as depicted in Fig. 3 (convergent views). We note that in the figure, for easier readability, we only show 3 days instead of 5. Interestingly, the machine also deem the water vapour as an important precursor to heatwaves in Indochina – more specifically, water vapor in India. Indeed, in this case the human view and the machine view are divergent, as depicted in Fig. 3 (divergent views). In the figure, we do not report the other two variable for the sake of readability.
@@
A way of seeing the simple experiment presented here is the following. For temperature, (human) domain knowledge luo2018synoptic partially overlaps the machine view, and they seem to provide a convergent (albeit not fully) view on the problem. For water vapour, (human) domain knowledge luo2018synoptic does not overlap the machine view, and provide a divergent view on the problem. The convergent view may reassure us that the results are trustworthy; yet, the divergent view can promote further scientific investigation that may lead to new knowledge on the topic. The convergent view that leads to trustworthiness is particularly relevant for critical applications in the applied sciences, such as in medicine, where wrong decisions may have catastrophic consequences. The divergent view that leads to further scientific investigation is instead relevant for both the natural and applied sciences, especially in those fields and processes where we have limited knowledge.
@@
Obviously, the relatively simple example shown here has some limitations. For instance, the variables used are limited and we are not including a set of data configurations that may be useful to further understand the problem at hand, such as changing the domain size. These limitations showcase the challenges that scientific XAI can encounter, and that needs to be tackled for this endeavour to have a chance of succeeding.
Learning through XAI, really?
@@
The perspective we present here was inspired by several authors that published on the topic of AI for science in the past few years, but perhaps one contribution stands out: the inspiring New York Times editorial by Steven Strogatz covering the winning of AlphaZero against Stockfish strogatz2018one . In that piece, Strogatz states: “What is frustrating about machine learning, however, is that the algorithms can’t articulate what they’re thinking. We don’t know why they work, so we don’t know if they can be trusted. AlphaZero gives every appearance of having discovered some important principles about chess, but it can’t share that understanding with us.”. He additionally cites Garry Kasparov (the former world chess champion) that stated: “we would say that its [AlphaZero] style reflects the truth. This superior understanding allowed it to outclass the world’s top traditional program despite calculating far fewer positions per second.” kasparov2018chess .
@@
While science is a rather different application than Go, we believe that leveraging explanation for scientific inquiry that may ultimately lead to knowledge discovery may not be a foolish path.
We should also add that some views suggesting that generative AI, including ChatGPT, and foundation models can autonomously generate new science may not be a plausible path, unless complex human reasoning is kept in the loop, at least until when we will achive artificial general intelligence. XAI is a possible path forward to keep complex human reasoning in the loop while bridging current AI power in discovery patterns and relationships in data.
XAI may also alleviate some of the risks that we may face when using AI for scientific discovery, that we share with Messeri and Crockett messeri2024artificial .

### v2

@@ Abstract
The framing of the problem is as follows. We want our AI approach to predict heatwaves in the Indochina peninsula 1 day ahead of time, using the prior 5 days. The associated learning task for our AI framework (based of a transformer neural network) is to correctly classify whether 1 day in the future corresponds or not to the onset day of a heatwave. We train our AI framework on time series data corresponding to 4 different variables that are relevant to heatwaves for the 5 days prior to the onset (or no onset) of a heatwave, and on a geographical domain larger than the Indochina peninsula. The training dataset is composed of 365 samples, the validation set of 77 samples, and the testing set also of 77 samples. The classification results show an overall accuracy on the test set of 88.3% (68/77), with a true positive (heatwave event) rate of 76.5% (13/18) and a true negative (no event) rate of 91.6% (55/59). From these results, we may argue that the machine learnt some ‘principles’ (statistical patterns in the data used) related to heatwaves in Indochina, and satisfies the assumption that we set in the previous section. We also note that, a transformer neural network may provide a causal model nichani2024transformers , and many works are devoted to this topic – see e.g., melnychuk2022causal ; rohekar2024causal .
@@
We would like now to understand what data the machine deemed important to learn these ‘principles’. To this end, we adopt post-hoc interpretability, and in particular the methods and evaluation framework presented in turbe2023evaluation , to guarantee that the relevance maps (i.e., the machine view) are correctly identified. We then compute the average machine view across all samples that are correctly classified, for the 4 variables considered, and for the 5 days prior to a heatwave striking that were used as inputs. We repeat the experiment a number of times, obtaining results that are similar to each other.
The first question we should ask ourselves is whether the ‘machine view’ just retrieved satisfies the pillars or requirements of XAI for science set out in the previous section. By evaluating the accuracy of the relevance maps, we attempt to satisfy the first requirement, namely accuracy. The possible inherent causality of the transformer model satisfies the second requirement, that we deem appropriate for this application. The reproducibility of the results for different repetitions of the experiment satisfies the third, namely reproducibility. Finally, the variables used are understandable to domain experts (i.e., weather and climate scientists), satisfying the fourth requirement. It is clear that further work should be undertaken to guarantee that the above requirements are indeed satisfied. But, for the sake of this thought experiment, we shall assume that this is the case.
@@
What can we learn from the ‘machine view’ just retrieved? We can look at what domain experts understand of heatwaves in Indochina. To this end, we found only one relevant reference luo2018synoptic , that shows how temperature anomalies in South China seem to be systematic precursors of Indochina’s heatwaves, and constitutes the human view on this problem (as reported in Fig. 3). Remarkably, this human view and the machine view are convergent, at least partially – they deem the same temperature data as important for heatwaves in Indochina in the days preceding these events, as depicted in Fig. 3 (convergent views). We note that in the figure, for easier readability, we only show 3 days instead of 5. Interestingly, the machine also deem the water vapour as an important precursor to heatwaves in Indochina – more specifically, water vapor in India. Indeed, in this case the human view and the machine view are divergent, as depicted in Fig. 3 (divergent views). In the figure, we do not report the other two variables for the sake of readability.
@@
A way of seeing the simple experiment presented here is the following. For temperature, (human) domain knowledge luo2018synoptic partially overlaps the machine view, and they seem to provide a convergent (albeit not fully) view on the problem. For water vapour, (human) domain knowledge luo2018synoptic does not overlap the machine view, and provide a divergent view on the problem (we shall note that we are working with domain experts on the specific example presented here).
The divergent view can promote further scientific investigation that may lead to new knowledge on the topic, and it is relevant for both the natural and applied sciences, especially in those fields and processes where we have limited knowledge. The convergent view may reassure us that the results are trustworthy, and it is particularly relevant for critical applications in the applied sciences, such as in medicine, where wrong decisions may have catastrophic consequences.
@@
Obviously, the relatively simple example shown here has some limitations. For instance, the accuracy of the machine view may be further tested, with e.g., other methods, the causality of the model should be further explored, and variables as well as the data configurations used should be expanded, to further assess the reproducibility of the problem and improve its understandability. These limitations showcase the challenges that scientific XAI can encounter, and that needs to be tackled for this endeavour to have a chance of succeeding. We shall discuss this aspect next.
Limitations and future outlook
From the example just introduced, it is clear that scientific XAI has some limitations and possible pitfalls, especially given the current approaches to XAI. I shall discuss them in the following, using the four requirements set for scientific XAI as a roadmap, and providing a final comprehensive future outlook on the subject.
Let us start from the accuracy requirement. To obtain what we called the machine view there exist several methods, grouped into two different categories, namely post-hoc and intrinsic (also referred to as ante-hoc) interpretability – see e.g., graziani2023global for a comprehensive review and taxonomy of interpretable AI. Post-hoc interpretability suffers from inaccuracy, aspect that may render futile the effort by domain expert to explain why certain data was used, if the data pinpointed is wrong. To this end, accuracy evaluation methods are as important, see e.g., turbe2023evaluation . Intrinsic interpretability may address the accuracy issues that plague post-hoc interpretability, yet they may be model-specific, human-constrained, and unable to provide machine views for complex deployed models that may not be easily accessible. Yet, several promising works are being carried out in the context of intrinsic interpretability – see e.g., alvarez2018towards ; zhang2022protgnn ; turbe2024protos . Another recent promising area is (intrinsic) compositional interpretability tull2024towards , that may help tie together different intrinsic interpretability viewpoints under a unified framework.
In terms of the causality requirement, the patterns and relationships identified by the machine may be just spurious correlations, without any causality, and may fool scientists towards chasing wrong answers to the ‘why’ question. Understanding whether causality is appropriate to the problem being investigated, and assessing whether the machine view is causal is a key aspect that scientists should take into account in answering ‘why’ the machine deemed important certain data. Indeed, several works are now exploring the intersection between XAI and causality carloni2023role , two fields that in the computer science community have drifted apart over the years beckers2022causal , but that should be kept together, especially in the natural and applied sciences. To this end, some promising works are emerging – see e.g., scholkopf2022causality ; janzing2020feature ; xu2020causality . In addition, Peters et al. book peters2017elements provides an introduction to causality for learning algorithms, while Pearl’s book offers an accessible roadmap and perspective on causal inference from both a statistical and a philosophical perspective pearl2018book .
In terms of reproducibility, we should be able to retrieve the same machine view repeatedly, for a given task and set of samples. In this context, some issues derive from the possible non-uniqueness of the machine view that interpretability methods may provide (especially post-hoc interpretability methods). This can hinder our understanding of why those machine views were used. To this end, answering the question ‘how’ was the machine view (i.e., the data deemed important) used by the machine can provide a more useful insight, as well pointed out by Rudin rudin2019stop .
Finally, in terms of understandability, the machine view should allow scientists and domain expert to tap into their body of knowledge, to allow for comparing machine and human views on a certain phenomenon of interest. This is partly connected to what discussed by Freiesleben et al. freiesleben2022scientific . More specifically, the model should be able to address a question understandable to scientists, with a machine view that can also be understood by scientists within their framework of domain-specific knowledge. The choice of data and data configuration is therefore critical, and frequently arbitrary, leaving a degree of subjectivity and arbitrariness to the problem. For instance, if we consider the growing field of domain adaptability (i.e., using an AI model for a different domain than the one it was trained on) long2015learning , the machine view may not be understandable to scientists. These issues are less pronounced in the context of fine-tuning a certain model, as the pre-training is usually accomplished within the same domain as the fine-tuning, hence understandable to domain experts working in the specific field.
The challenges just presented, together with the overall premise of being able to transfer knowledge to human scientists via the machine view we introduced, may let scientific XAI seem like a futile endeavour. Yet, we believe that there is concrete hope. Let me take a brief detour, and explain why.
@@
The perspective presented here is built on several works that have been published in the past few years within the context of AI, but perhaps one contribution stands out: the inspiring New York Times editorial by Steven Strogatz covering the winning of AlphaZero against Stockfish strogatz2018one . In that piece, Strogatz states: “What is frustrating about machine learning, however, is that the algorithms can’t articulate what they’re thinking. We don’t know why they work, so we don’t know if they can be trusted. AlphaZero gives every appearance of having discovered some important principles about chess, but it can’t share that understanding with us.”. He additionally cites Garry Kasparov (the former world chess champion) that stated: “we would say that its [AlphaZero] style reflects the truth. This superior understanding allowed it to outclass the world’s top traditional program despite calculating far fewer positions per second.” kasparov2018chess .
@@
While science is a rather different application than Go, we believe that leveraging explanation for scientific inquiry that may ultimately lead to knowledge discovery may not be a foolish path. XAI is a possible path forward to keep complex human reasoning in the loop while bridging current AI power in discovering patterns and relationships in data. XAI may also alleviate some of the risks that we may face when using AI for scientific discovery, that we share with Messeri and Crockett messeri2024artificial .
We add that some views suggesting that current generative AI, including ChatGPT, and foundation models can autonomously generate new science may not be a plausible path unless complex human reasoning (provided by human scientists and domain experts) is kept in the loop, at least until when we will achieve artificial general intelligence Superintelligence .